# On the Interchangeability of Positional Embeddings in Multilingual Neural Machine Translation Models

### v1

@@ 5.1 Swapping out Positional Embeddings
Tables 4 and 5 demonstrate the performance of several experimental setups that involve swapping of Positional Embeddings and followed by fine-tuning (if applicable).
Performance Degradation Post-Swap
As expected, a significant drop in performance is observed when Positional Embeddings are removed, compared to the based, with further deterioration occurring if the new type of Positional Embeddings are directly swapped in without subsequent fine-tuning. However, a key observation in this scenario is that performance does not completely collapse, suggesting that positional information may be partially encoded within the model itself. This indicates that attributing the entirety of positional information to the Positional Embeddings alone may not be entirely accurate.
Fine-tuning is key to performance recovery
@@
When Positional Embeddings such as RoPE or ALiBi are introduced, across various fine-tuning setups, we consistently observe that the performance can be regained through fine-tuning. This demonstrates that it is feasible to modify Positional Embeddings at later stages and still retain the performance of the base model.
Efficacy of the Positional Embedding switch
@@
A notable finding in this setting is that fine-tuning a model originally using sinusoidal embeddings yielded performance comparable to models fine-tuned after the switch to using RoPE or ALiBi embeddings. This further reinforces the conclusion that Positional Embeddings in pre-trained transformer models can be safely tweaked even in later stages. Although sinusoidal, rotary, and ALiBi embeddings show similar performance in long-context document-level tests used in the context of this study. The advantage of switching to RoPE or ALiBi is that further length extrapolation may now be feasible in these transformer models, as demonstrated by the findings of Su et al. (2024) and Press et al. (2022).
NoPE is not a solution in Encoder-Decoder models
Another key observation in our setting is that the findings of Kazemnejad et al. (2023) about explicit Positional Embeddings not being necessary does not generalize to Encoder-Decoder models. As we observe that no fine-tuning strategy is capable of bridging the gap and regaining performance.
Minimalistic Fine-tuning is sufficient
@@
Across all the fine-tuning strategies evaluated, minimalistic LoRA fine-tuning, as well as FFT and LoRA, produced comparable performance. This suggests that extensive parameter adjustments are not necessary; adding LoRA parameters solely to the self-attention modules, where the relative Positional Embedding modules are integrated, is sufficient for adaptation. This approach is highly parameter-efficient, requiring only 2.35M trainable parameters (1.11%).
Base SINE NoPE RoPE ALiBi
IT2 N R A FFT L ML FFT L ML FFT L ML FFT L ML
Flores Indic-En 59.5 34.9 31.5 23.3 59.6 59.2 59.3 42.8 41.7 41 59.4 59.0 59.0 59.3 58.7 58.7
IN22-Conv Indic-En 55.8 38.6 38.0 35.0 55.8 55.6 55.6 44.8 44.3 43.7 55.6 55.3 55.4 55.3 55.1 55.2
IN22-Gen Indic-En 62.0 34.1 28.0 19.5 61.9 61.6 61.6 42.5 41.1 40.3 61.6 61.4 61.3 60.7 60.9 60.8
Flores En-Indic 47.9 31.5 27.0 15.9 48.1 47.7 47.5 39.0 37.6 36.6 48.2 47.7 47.5 49.4 47.6 47.5
IN22-Conv En-Indic 44.5 35.2 31.3 26.1 42.4 42.0 42.5 38.7 37.5 37.2 42.5 42.0 42.0 42.8 42.0 42.1
IN22-Gen En-Indic 48.3 30.1 24.9 13.4 46.3 46.1 46.7 36.9 35.1 33.7 46.3 46.2 46.0 46.4 46.0 45.9
Table 4: ChrF++ scores obtained when swapping out the Positional Embeddings and fine-tuning, evaluated across Sentence-level benchmarks. N denotes a direct transition (without any tuning) to No-Positional Embeddings, R denotes a direct transition to Rotary Positional Embeddings, and A denotes a direct transition to ALiBi Positional Embeddings. “FFT” represents Full Fine-Tuning, “L” signifies LoRA fine-tuning, and “ML” refers to minimalistic LoRA fine-tuning (applied exclusively to the Self-Attention block).
Base SINE NoPE RoPE ALiBi
IT2 N R A FFT L ML FFT L ML FFT L ML FFT L ML
Flores Indic-En 60.9 28.6 12.4 5.6 61.3 61.1 61.2 37.8 34.9 33.6 61.0 60.8 60.9 59.3 59.2 59
IN22-Conv Indic-En 27.5 10.2 2.3 1.1 32.7 32.0 31.5 15.1 13.4 13.7 33.0 32.8 33.0 30.2 29.7 28.1
ALT Indic-En 65.2 39.5 35.4 26.8 65.2 65.2 64.8 49.2 48.0 47.1 65.0 65.0 64.9 64.7 64.8 64.8
Flores En-Indic 49.5 23.8 14.8 4.8 49.8 49.2 49.1 33.1 29.2 26.8 49.9 49.1 49.1 49.4 48.6 48.3
IN22-Conv En-Indic 25.4 7.7 2.7 1.3 24.5 23.4 22.9 11.1 9.5 8.0 24.1 23.5 22.6 23.3 20.8 21.1
ALT En-Indic 54.9 36.7 31.9 19.4 55.3 55.0 54.9 45.6 43.9 43.1 55.3 55.0 55.0 55.4 55.0 55.0
Table 5: ChrF++ scores obtained when swapping out the Positional Embeddings and or fine-tuning, evaluated across Document-level benchmarks averaged across languages. N denotes a direct transition (without any tuning) to No-Positional Embeddings, R denotes a direct transition to Rotary Positional Embeddings, and A denotes a direct transition to ALiBi Positional Embeddings. “FFT” represents Full Fine-Tuning, “L” signifies LoRA fine-tuning, and “ML” refers to minimalistic LoRA fine-tuning (applied exclusively to the Self-Attention block).
NoPE RoPE ALiBi
Flores Indic-En 38.8 59.4 58.4
IN22-Conv Indic-En 14.0 33.2 31.6
ALT Indic-En 52.5 64.2 63.9
Flores En-Indic 37.2 49.3 49.1
IN22-Conv En-Indic 17.3 26.0 25.8
ALT En-Indic 48.4 55.2 55.2
Table 6: ChrF++ scores of models trained from scratch with various Positional Embeddings evaluated across Document-Level benchmarks averaged across languages.
NoPE RoPE ALiBi
Flores Indic-En 44.9 57.9 57.5
IN22-Conv Indic-En 45.5 54.2 54.0
IN22-Gen En-Indic 40.9 48.1 47.9
Flores En-Indic 40.6 47.6 47.6
IN22-Conv En-Indic 41.0 44.4 44.2
IN22-Gen Indic-En 44.9 60.3 59.9
Table 7: ChrF++ scores of models trained from scratch with various Positional Embeddings evaluated across Sentence-Level benchmarks averaged across languages.
5.2 Comparison with Training from scratch
Tables 6 and 7 demonstrate the performance trends of the models with the respective Positional Embedding trained from scratch. We compare these with the results in Tables 4 and 5. The comparison shows that models with swapped Positional Embeddings either outperform or are on par with their counterparts trained from scratch. However, training from scratch is resource-intensive, whereas the proposed approach of post-hoc Positional Embedding swapping proves to be a more efficient alternative in terms of both computation and data. Additionally, models trained from scratch using NoPE slightly outperform their swap counterparts, but remain suboptimal compared to any other Positional Embeddings (SINE / ALiBi / RoPE), further reinforcing the prior finding that NoPE is not a viable option for Encoder-Decoder models.
@@ 6 Conclusion
In this study, we have demonstrated that sinusoidal Positional Embeddings can be efficiently replaced with RoPE or ALiBi in a standard transformer model. We show that this substitution can be achieved in a parameter-efficient manner by simply introducing additional parameters within the self-attention module. Our findings also indicate that NoPE is not a viable solution for Encoder-Decoder models, underscoring the necessity of explicit positional information in such architectures. Although our experiments are focused on the machine translation (MT) task, we believe that this approach can be extended to other Encoder-Decoder models or Encoder-only models to enhance their long-context capabilities.
@@ 7 Limitations
  • Our investigation does not extend to examining the interchangeability between two relative PEs or from relative to sinusoidal PEs. This is due to the fact that older models predominantly employed sinusoidal PEs, whereas contemporary SOTA LLMs tend to favor RoPE, which has demonstrated robustness in handling long contexts. Therefore, a practically viable setup would involve a switch from sinusoidal to relative and not the other way around.
  • Our study concludes with demonstrating that interchangeability is feasible. Further, we also demonstrate that fine-tuning with sinusoidal embeddings also results in nearly similar performance. However, it is important to note that prior works demonstrate that RoPE provides superior length generalization capabilities, however as we do not have any long-context benchmark at our disposal as of now, we leave this aspect to future work.

### v2

@@ 5 Results and Discussions
Figure 2: ChrF++ scores on the Document-level benchmarks of baselines involving 3 fine-tuning setups FFT, LORA, min-LORA comparing 3 types of postional embeddings namely SINE, ALIBI, ROPE along with the pre-trained IT2 performance baseline.
Figure 3: ChrF++ scores on the Sentence-level benchmarks of baselines involving 3 fine-tuning setups FFT, LORA, min-LORA comparing 3 types of postional embeddings namely SINE, ALIBI, ROPE along with the pre-trained IT2 performance baseline.
@@ 5.1 Swapping out Positional Embeddings
Figures 3 and 2 demonstrate the performance of several experimental setups that involve swapping of Positional Embeddings and followed by fine-tuning (if applicable). Detailed results are also presented in Tables 5 and 6.
Fine-tuning is key to performance recovery.
@@
When Positional Embeddings such as ROPE or ALIBI are introduced, we observe a significant drop in performance as the model is not trained to work with these. However, across various fine-tuning setups, we consistently observe that the performance can be regained through fine-tuning (see Figure 3). This demonstrates that it is feasible to modify Positional Embeddings at later stages and still retain the performance of the base model.
Efficacy of the Positional Embedding switch.
@@
A notable finding in this setting is that fine-tuning a model originally using sinusoidal embeddings yielded performance comparable to models fine-tuned after the switch to using ROPE or ALIBI embeddings on sentence-level benchmarks as observed in Figure 3. This further reinforces the conclusion that Positional Embeddings in pre-trained transformer models can be safely tweaked even in later stages. Although sinusoidal, rotary, and ALIBI embeddings show similar performance in sentence-level test sets. The advantage of switching to ROPE or ALIBI can be observed in case of document-level test sets in line with the findings of Su et al. (2024) and Press et al. (2022).
Figure 4: Implicit MQM scores and normalized formula-based MQM scores on the IN22-Conv benchmark, averaged across Top-12 languages highlighted in Table 3
Cross-lingual length-generalization.
Our fine-tuning dataset primarily comprised sentence-level data, as detailed in Section 3.2.1, with only a limited amount of long-context or paragraph-level data available for four Dravidian languages accounting for less than 0.05% of the complete fine-tuning data mixture (see Table 1). The token distribution across this dataset is illustrated in Table 2. While most languages lacked long-context data, evaluation on document-level benchmarks still demonstrated strong performance on these, as seen in Figure 2. This indicates that fine-tuning with a small amount of long-context data introducing relative positional embeddings can enable robust cross-lingual length generalization. This finding is particularly important because long-context data is often scarce across many languages. Nevertheless, effective adaptation to longer contexts was achieved through cross-lingual generalization, even with minimal long-context data in select languages.
Minimalistic LORA is on par with LORA
@@
Across all the fine-tuning strategies evaluated, minimalistic LORA (min-LORA) fine-tuning, as well as FFT and LORA, produced comparable performance on sentence-level benchmarks. This suggests that extensive parameter adjustments are not necessary; adding LORA parameters solely to the self-attention modules, where the relative Positional Embedding modules are integrated, is sufficient for adaptation. This approach is highly parameter-efficient, requiring only 2.35M trainable parameters (1.11%). On the document-level benchmarks, we observe that both LORA and min-LORA are on par suggesting that min-LORA is adequate. However these parameter-efficient approaches slightly lag behind the respective FFT baseline.
Throughput
ROPE has lower throughput compared to Sinusoidal embeddings due to its higher number of operations. Since the rotation function in ROPE is applied at every layer, it has a time complexity of $\mathcal{O}(n)$, whereas Sinusoidal embeddings are added only once, after the token embeddings, resulting in $\mathcal{O}(1)$ complexity. The exact decoding rates (in tokens per second) are illustrated in Figure Figure 5. An efficient implementation of ALIBI Press et al. (2022) achieves $\mathcal{O}(1)$ complexity, making it slightly slower than Sinusoidal but still significantly faster than ROPE. Therefore, for a basic sentence-level MT system, Sinusoidal PEs offer the best efficiency in terms of throughput and performance. However, for document-level or longer contexts, ROPE provides notable qualitative advantages, although it comes with a decrease in throughput.
Figure 5: Throughput in terms of token/sec. Higher is better.
5.2 Qualitative Analysis
In the preceding section, we conducted a string-metrics-based evaluation on standard benchmarks such as IN22-Conv, ALT, and FLORES, covering both sentence-level and document-level variants. However, the effectiveness of these string-based metrics in evaluating translation quality, especially in longer-context scenarios where factors like semantic coherence, coreference resolution, and continuity are critical, remains underexplored.
Therefore, for a more fine-grained analysis we conduct GPT-4o based evaluation as described in Section 3. We focus exclusively on the FFT variants for GPT-4o evaluation, as they were the highest-performing systems. Figure 4 demonstrates two types of evaluations: an implicit score assigned by GPT-4o and a calculated score based on MQM annotations generated by GPT-4o. Both metrics clearly indicate that ROPE-based baselines outperform the SINE and ALIBI counterparts, with ALIBI proving superior to the SINE baseline. We also observe that, in a relative sense, the rankings derived from the implicit scores generally align with those based on the calculated scores. These results underscore the effectiveness of the proposed approach, showing that substituting positional embeddings had a notable positive impact on tasks involving longer contexts. We also present a table of errors identified by GPT-4o on the same sample conversation across the three systems being compared: ROPE, SINE, and ALIBI in Tables 7, 8 and 9 respectively. The MQM annotations clearly show that the translations produced by the ROPE model exhibit less severe errors, followed by ALIBI. In contrast, the SINE model has a higher number of errors with greater severity. This aligns with the earlier observations, confirming that the ROPE models are indeed superior to the SINE and ALIBI counterparts.
@@ 6 Conclusion
In this study, we have demonstrated that sinusoidal Positional Embeddings can be efficiently replaced with ROPE or ALIBI in a standard transformer model. We show that this substitution can be achieved in a parameter-efficient manner by simply introducing additional parameters within the self-attention module. Although our experiments are focused on the machine translation (MT) task, we believe that this approach can be extended to other Encoder-Decoder models or Encoder-only models to enhance their long-context capabilities.
@@ 7 Limitations
  • We use string-based automatic evaluation and GPT-4o as an evaluator, but human evaluation may be needed to further validate the results and reinforce our findings.
  • The benchmarks we test in this study are smaller in size due to the limited availability of long-context MT datasets for Indian languages.
@@ 8 Acknowledgements
Appendix A Appendix
A.1 Scores
The raw scores corresponding to fig. 2 and fig. 3 shown in Table table 5 and table 6.
FT LoRA min LoRA
DATASET IT2 SINE ALIBI ROPE SINE ALIBI ROPE SINE ALIBI ROPE
IN22 Gen 48.4 47.7 47.6 47.8 47.0 47.1 46.9 47.1 46.9 47.1
IN22 Conv 44.6 43.6 43.8 43.8 42.9 43.1 42.9 43.0 43.1 43.1
En-Indic FLORES 47.8 48.3 48.3 48.2 47.8 47.7 47.7 47.7 47.6 47.6
IN22 Gen 62.0 62.0 61.2 61.8 61.6 61.0 61.4 61.6 60.9 61.5
IN22 Conv 55.7 55.3 55.0 55.1 55.4 54.7 55.3 55.5 55.2 55.4
Indic-En FLORES 59.5 59.7 59.2 59.4 59.4 58.8 59.2 59.3 58.8 59.1
Table 5: ChrF++ scores on sentence-level evaluation benchmarks.
FT LoRA min LoRA
DATASET IT2 SINE ALIBI ROPE SINE ALIBI ROPE SINE ALIBI ROPE
ALT 55.0 55.2 55.4 55.3 55.1 55.0 55.0 55.0 55.0 54.9
IN22 Conv 32.4 36.2 42.4 45.1 29.6 33.5 38.6 29.3 34.1 39.4
En-Indic FLORES 49.7 50.3 49.9 50.2 49.6 48.9 49.3 49.3 48.5 49.2
ALT 65.2 65.2 64.8 65.1 65.2 64.9 65.0 65.2 64.9 64.9
IN22 Conv 32.4 40.3 51.8 56.6 38.8 47.1 55.1 35.3 43.8 53.7
Indic-En FLORES 61.2 61.2 60.0 61.2 61.1 59.4 60.9 61.2 59.1 60.9
Table 6: ChrF++ scores on document-level evaluation benchmarks.
A.2 Qualitative Examples
The MQM annotations generated by GPT-4 for a test conversation across for each baseline is presented in Tables 7, 8 and 9.
Type Category Classification Span Text Explanation
Fluency Awkward Minor please sit down Slightly awkward phrasing. ‘Please have your seat’ is more natural.
Fluency Inconsistency Minor on six months, on three months, or every month Inconsistent phrasing in intervals. Use ‘half-yearly, quarterly or monthly intervals.
Accuracy Mistranslation Major And I also know its benefits. Incorrectly translated. Should be ‘I want to know the benefits.
Accuracy Omission Major You can opt for that. Omitted in the translation.
Accuracy Mistranslation Major Please write my phone number! Redundant repetition at the end which was not present in the source.
Accuracy Mistranslation Critical after the completion of the first five years of the policy Misleading regarding the conditions of the death benefit. Should align with ‘On death after completion of five policy years but before the date of maturity…
Table 7: ROPE Translation Errors with Classification
Type Category Classification Span Text Explanation
Fluency Grammar Minor The entire information about the process Awkward sentence structure. ‘The information about the process can be’ would be more fluent.
Accuracy Mistranslation Major Your name and age are Devang Saikia. I am 39 years old This part is translated inaccurately. The source asks for name and age, but the translation incorrectly states them directly.
Accuracy Omission Major You can opt for that. Translation omits ‘What is the minimum basic sum assured in this policy?’ and other questions related to policy details.
Accuracy Addition Major What is the minimum age? What is the minimum age? What is the minimum age? Text is repeated unnecessarily, creating confusion.
Accuracy Mistranslation Major You will have to surrender the policy regularly Mistranslation; the context is about having completed certain years, not regular surrender.
Accuracy Mistranslation Critical premium annually on the death of the policy Incorrect context; it should refer to payment intervals, not related to death of the policy.
Table 8: ALIBI Translation Errors with Classification
Type Category Classification Span Text Explanation
Accuracy Mistranslation Critical Thank you for giving my name and age. Incorrect translation. Should be ‘Thank you for letting us know your name and age.
Accuracy Mistranslation Critical What will be the premiums paid during the first five years of the policy? Incorrect translation. Should be ‘On death during the first five years, sum assured on death shall be payable.
Accuracy Mistranslation Critical But do I want to know the full amount of the policy before the end of the year? Incorrect translation. Should be ‘Can I surrender the policy before the maturity date? What will be the surrender value?
Accuracy Mistranslation Critical what will be the premiums paid before the end of the policy? Incorrect translation. Should refer to death benefits, not about premiums.
Accuracy Mistranslation Critical I can give you the information on the phone. Text unrelated to the source and incorrectly translated.
Accuracy Mistranslation Critical If it is less than three months old, then it will be enough to pay the premiums. Incorrect translation. Policy is misunderstood. Should reflect surrender conditions.
Table 9: SINE Translation Errors with Classification